# Motion Synthesis with Sparse and Flexible Keyjoint Control

### v1

@@ 2 Related Works
Recent advances in data-driven human motion synthesis have enabled the generation of highly realistic and diverse human motions. Modern generative modeling techniques—such as generative adversarial networks (GANs) [65], variational autoencoders (VAEs) [16], token-based models [18], and diffusion models [58]—have successfully captured the complexity of human motion dynamics. Many works focused on synthesizing motions that adapt to high-level semantic conditions (e.g., action categories [15, 45, 58, 8, 23], text descriptions [17, 18, 70, 69, 16, 58, 72, 46, 40, 47], audio [33, 14, 1, 53, 59, 38, 49]) and surrounding environments (e.g., interacting objects [66, 13, 37, 75, 6, 32, 57, 4, 36] or scenes [26, 62, 25, 39, 34, 76, 68, 7, 24, 63, 56]). These methods bridge the gap between user intent and complex motion generation, enabling intuitive applications for animation.
@@
In addition to high-quality motion generation from high-level intents (e.g., text prompts), allowing spatially precise control is critical for practical applications, particularly in interactive animation workflows. A foundational task in this domain is motion in-betweening [3, 5, 31, 19, 73, 20, 20, 12, 50, 43, 30], which synthesizes plausible transitions between keyframes defined as full-body information. Early works explored interpolation using splines such as Bézier curves [51, 42, 35], but these methods often lacked the flexibility to model complex human motion. Recent neural approaches have improved robustness: Qin et al. [50] proposed a two-stage deterministic network for motion completion, while CondMDI [9] leveraged generative capabilities of diffusion models and extended their abilities to incorporate text instructions.
@@
Beyond keyframe interpolation, recent methods focus on joint-level spatial control for advanced motion synthesis. PriorMDM [52] pioneered joint-aware generation by fine-tuning a pre-trained motion diffusion model MDM [58]. GMD [52] introduces classifier-free guidance to diffusion models to enable root joint trajectory control, and OmniControl [64] generalized the control capability to arbitrary joints via ControlNet [71] modules with spatial guidance at inference. MotionLCM [10] further incorporated ControlNet into a motion latent space with a consistency model. Other approaches explored latent space optimization: TLControl [60] trained part-based discrete latent codes and performed inference-time code optimization, while ControlMM [48] exploited spatial guidance into a masked motion model via token optimization. DNO [27] optimized diffusion noise via a differentiable objective function defined through global joint position. A recent concurrent work, Studer et al. [55] introduced a factorized diffusion approach with Bézier curve parameterization, though it omits text conditioning.
Figure 2: From control input, we first synthesize the keyjoint trajectory $\mathbf{C}$ as an intermediate representation. For explicit control, the signal $\mathbf{s}$ is combined with the binary indicator $\mathbf{m_{s}}$. We impute $\mathbf{m_{s}}\circ\mathbf{s}$ to the keyjoint trajectory and denoise it through keyjoint trajectory model $\mathcal{D}_{\mathbf{c}}$ at each diffusion step. For implicit control, an objective function $\mathcal{F(\cdot)}$ is defined over the keyjoint $\mathcal{J}$ and optimized thorough diffusion latent optimization. The resulting keyjoint trajectory $\mathbf{C}$ is then imputed into the full-body motion completion model $\mathcal{D}_{\mathbf{x}}$ to synthesize the final full-body motion $\mathbf{X}$.
@@ 3.1.1 Motion Generation with Explicit Keyjoint Control
During training, the denoising network incorporates an imputation strategy with a randomly generated mask, earning the flexibility to handle arbitrary control signals. We select a timestep $t$ from the uniform distribution over $[1,T]$ and generate a random spatio-temporal mask $\mathbf{m_{C}}\in\{0,1\}^{N\times d}$. This mask $\mathbf{m_{C}}$ indicates which joints will be used for conditional motion synthesis. We then impute the noisy sample $\mathbf{C}_{t}$ with the corresponding values from the clean sample $\mathbf{C}_{0}$:
@@ 3.1.2 Motion Generation with Implicit Keyjoint Control
Our functional constraint can allow trajectory input without exact per-frame timestamps, namely time-agnoistic motion control. One critical limitation of explicit spatial motion control is the requirement for the exact frame-wise timing. Such information is highly challenging for an animator to accurately annotate, limiting their practicality in real-world scenarios.
@@ 4.1 Explicit Keyjoint Control
Table 1: Quantitative evaluation of the sparse joint control. We validate our framework under a highly sparse control setting, which uses only $0.454\%$ of total joint signal as control input. Bold represents the best value, and underlined represents the second-best.
@@ 4.2 Implicit Keyjoint Control
We compare our approach against both the original implementation of [27], which directly generates full-body motion, and our single-stage variant (referred to as “Ours w/o decomposed"). As presented in Table 3, our decomposed pipeline consistently improves constraint error while enhancing motion quality. By applying objectives at the keyjoint trajectory level rather than at the full-body motion synthesis stage, our approach better satisfies objective constraints while preserving the realism of the full-body motion after completion.

### v2

@@ 2 Related Works
Recent advances in data-driven human motion synthesis have enabled the generation of highly realistic and diverse human motions. Modern generative modeling techniques—such as generative adversarial networks (GANs) [72], variational autoencoders (VAEs) [17], token-based models [19], and diffusion models [65]—have successfully captured the complexity of human motion dynamics. Many works focused on synthesizing motions that adapt to high-level semantic conditions (e.g., action categories [16, 52, 65, 9, 26], text descriptions [18, 19, 77, 76, 5, 17, 65, 79, 53, 45, 54, 20], audio [38, 15, 1, 60, 66, 43, 56]) and surrounding environments (e.g., interacting objects [73, 14, 42, 82, 7, 37, 64, 4, 41] or scenes [31, 69, 30, 44, 39, 83, 75, 8, 27, 70, 63, 28]). These methods bridge the gap between user intent and complex motion generation, enabling intuitive applications for animation.
@@
In addition to high-quality motion generation from high-level intents (e.g., text prompts), allowing spatially precise control is critical for practical applications, particularly in interactive animation workflows. A foundational task in this domain is motion in-betweening [3, 6, 36, 21, 80, 22, 22, 13, 57, 50, 35, 29], which synthesizes plausible transitions between keyframes defined as full-body information. Early works explored interpolation using splines such as Bézier curves [58, 48, 40], but these methods often lacked the flexibility to model complex human motion. Recent neural approaches have improved robustness: Qin et al. [57] proposed a two-stage deterministic network for motion completion, while CondMDI [10] leveraged generative capabilities of diffusion models and extended their abilities to incorporate text instructions.
@@
Beyond keyframe interpolation, recent methods focus on joint-level spatial control for advanced motion synthesis. PriorMDM [59] pioneered joint-aware generation by fine-tuning a pre-trained motion diffusion model MDM [65]. GMD [59] introduces classifier-free guidance to diffusion models to enable root joint trajectory control, and OmniControl [71] generalized the control capability to arbitrary joints via ControlNet [78] modules with spatial guidance at inference. MotionLCM [11] further incorporated ControlNet into a motion latent space with a consistency model. Other approaches explored latent space optimization: TLControl [67] trained part-based discrete latent codes and performed inference-time code optimization, while ControlMM [55] exploited spatial guidance into a masked motion model via token optimization. DNO [32] optimized diffusion noise via a differentiable objective function defined through global joint position. A recent concurrent work, Studer et al[62] introduced a factorized diffusion approach with Bézier curve parameterization, though it omits text conditioning.
Figure 2: From control input, we first synthesize the keyjoint trajectory $\mathbf{C}\in R^{N\times d}$ as an intermediate representation. For explicit control, the signal $\mathbf{s}$ is combined with the binary indicator $\mathbf{m_{s}}\in\{0,1\}^{N\times d}$. We impute $\mathbf{m_{s}}\circ\mathbf{s}$ to the keyjoint trajectory and denoise it through keyjoint trajectory model $\mathcal{D}_{\mathbf{c}}$ at each diffusion step. For implicit control, an objective function $\mathcal{F(\cdot)}$ is defined over the keyjoint $\mathcal{J}$ and optimized thorough diffusion latent optimization. The resulting keyjoint trajectory $\mathbf{C}$ is then imputed into the full-body motion completion model $\mathcal{D}_{\mathbf{x}}$ to synthesize the final full-body motion $\mathbf{X}\in R^{N\times D}$.
@@ 3.1.1 Motion Generation with Explicit Keyjoint Control
[…]
@@ 3.1.2 Motion Generation with Implicit Keyjoint Control
Our functional constraint can allow trajectory input without exact per-frame timestamps, namely time-agnostic motion control. One critical limitation of explicit spatial motion control is the requirement for the exact frame-wise timing. Such information is highly challenging for an animator to accurately annotate, limiting their practicality in real-world scenarios.
@@ 4.1 Explicit Keyjoint Control
[…]
@@ 4.2 Implicit Keyjoint Control
We compare our approach against both the original implementation of [32], which directly generates full-body motion, and our single-stage variant (referred to as “Ours w/o decomposed”). As presented in Table 3, our decomposed pipeline consistently improves constraint error while enhancing motion quality. By applying objectives at the keyjoint trajectory level rather than at the full-body motion synthesis stage, our approach better satisfies objective constraints while preserving the realism of the full-body motion after completion.
@@ 5 Conclusion
Acknowledgements
This work was supported by Creative-Pioneering Researchers Program through Seoul National University and BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2025. Inwoo Hwang is supported by Hyundai Motor Chung Mong-Koo Foundation.
Supplementary Material
In the supplementary materials, we describe the implementation details (Sec. 6) and present ablation results on different sampling strategy (Sec. 7). Please refer to the supplementary video on our project page for additional qualitative results.
6 Further Details
6.1 Model Details
We designed our each model using a diffusion framework based on DDPM [25], incorporating the U-Net architecture introduced by [33]. The training was conducted with the AdamW optimizer [47], where we set the learning rate to $1\times 10^{-4}$ and applied a weight decay of $1\times 10^{-2}$. For inference, we utilized classifier-free guidance with a weight $\mathbf{w}=2.5$. Additional details on the hyperparameters for both the network architecture and diffusion process are summarized in Table 8. Both models were trained using a diffusion process with 50 steps. The key-joint diffusion model is comparatively smaller in size in contrast to the full-body completion diffusion model, which consists of a channel dimension of 128.
Hyperparameter Key-Joint Model Full-Body Model
Learning rate 1e-4 1e-4
Optimizer Adam W Adam W
Weight decay 1e-2 1e-2
Batch size 64 64
Channels dim 128 512
Channel multipliers [2,2,2,2][2,2,2,2][ 2 , 2 , 2 , 2 ] [2,2,2,2][2,2,2,2][ 2 , 2 , 2 , 2 ]
Variance scheduler Cosine [49] Cosine [49]
Diffusion steps 50 50
Diffusion variance $\tilde{\beta}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ $\tilde{\beta}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$
EMA weight ($\beta$) 0.9999 0.9999
Guidance weight ($\mathbf{w}$) 2.5 2.5
Table 8: Hyperparameters of each model
6.2 Baseline Details
For the baselines OmniControl [71], TLControl [67], MotionLCM [11], and DNO [32], we utilize the officially released checkpoints for evaluation. For CondMDI [10], in order to support arbitrary joint-level control with their strategy, we train the model using global position representations for all joints.
6.3 Goal-Driven Motion Synthesize Task
[…]
We evaluate our approach to goal-driven motion synthesis task by training a unified network that integrates multiple scenarios. In each scenario, control signals are provided as an initial pose paired with target control joints at the final frame. Specifically, reaching target hand positions [2] focuses on controlling the right-hand position, climbing with rock constraints [74] involves controlling both hands and feet, and sitting with hand control [81] requires controlling both hands at the final frame. Our unified network is trained to address multiple dynamics and control settings simultaneously.
Dataset Description
We collect a variety of tasks that require control over different target joints at the final frame and involve multiple motion dynamics. For the scenario of reaching target hand positions, we utilize the dataset from [2]. Specifically, we extract sequences of reaching motions and augment them by mirroring the left-hand reaching motions to the right hand, thus generating a total of 3,138 right-hand reaching sequences. To define the goal position, we identify the farthest point reached by the hand from its initial location. In our experimental setup, 2,510 samples are designated for training. In the case of the climbing with rock constraints scenario, we leverage the dataset from [74]. We carefully select sequences that depict the subject detaching from one climbing rock and securely reaching for another. This process yields 156 motion samples. For the sitting with hand control scenario, we use the dataset from [81]. We extract 160 sequences that begin with the subject in a stable position and end when they are seated in a chair. To ensure consistency across all tasks, we standardize the dataset by aligning the subject’s face direction to the $+z$ axis at the initial frame and setting the root position at the origin.
7 Ablation Study on Sampling Strategy
We perform a quantitative evaluation of various sampling strategies for diffusion models. Specifically, we train a keyjoint trajectory model and a full-body completion model using 50-step diffusion models and apply both DDPM and DDIM-based diffusion sampling strategies. The results demonstrate that our method maintains consistent performance even with 5-step diffusion sampling and continues to outperform other baselines, achieving high precision and natural motion quality, as supported by Figure 3 of the main paper.
Sampling Frame Select FID $\downarrow$ Control Err. ($m$$\downarrow$ R-precision (Top-3) $\uparrow$ Div. $\rightarrow$ Foot Skating $\downarrow$
[…]
DDPM (50) $r=1$ 0.127 0.019 0.681 9.518 0.071
$r=2$ 0.128 0.019 0.680 9.539 0.070
$r=5$ 0.148 0.024 0.681 9.554 0.069
$r=10$ 0.171 0.027 0.678 9.402 0.074
$r=20$ 0.195 0.033 0.677 9.575 0.064
$r=30$ 0.224 0.036 0.673 9.674 0.061
$r=60$ 0.263 0.044 0.659 9.627 0.062
DDIM (10) $r=1$ 0.136 0.019 0.678 9.559 0.075
$r=2$ 0.141 0.020 0.681 9.574 0.074
$r=5$ 0.158 0.022 0.682 9.599 0.073
$r=10$ 0.186 0.025 0.675 9.632 0.069
$r=20$ 0.222 0.030 0.681 9.644 0.067
$r=30$ 0.254 0.037 0.673 9.621 0.063
$r=60$ 0.297 0.041 0.661 9.615 0.065
DDIM (5) $r=1$ 0.127 0.019 0.678 9.528 0.072
$r=2$ 0.140 0.019 0.686 9.573 0.073
$r=5$ 0.164 0.022 0.678 9.582 0.072
$r=10$ 0.202 0.025 0.674 9.543 0.065
$r=20$ 0.236 0.031 0.665 9.521 0.069
$r=30$ 0.338 0.037 0.655 9.356 0.065
$r=60$ 0.658 0.045 0.622 9.147 0.073
Table 9: Quantitative evaluation of various sampling strategies for diffusion models.
8 Additional Results
8.1 Results on Challenging or Unseen Scenarios
We further conduct experiments using manually specified control signals in challenging forms (e.g., S-curves or straight lines), which are unseen and difficult scenarios. These results highlight the robustness and generalization capability of our method (Figure 6).
Figure 6: Qualitative results on manually specified challenging control signals, such as S-curves and straight paths. Our method successfully follows the intended trajectories, demonstrating strong generalization to unseen and difficult motion constraints.
8.2 Results on More Expressive Prompts
We provide results for more expressive and complex prompts in Figure 7. Our method generates rich, detailed, and vivid motions, demonstrating its ability to capture expressive textual descriptions while faithfully adhering to sparse control signals such as “flying like an airplane,” “extending arms,” “walks backward,” and “walking up the stairs”.
Figure 7: Motion generation results for expressive prompts. Our method produces vivid and diverse motions in response to complex textual descriptions while respecting the given sparse control signals.